Calculate the volume (in mL) of a 26.0% (m/V) solution containing 14.4 g NaCl.
Answer #1

Answer:

Step 1:

Explanation:

Percent (m/V) is the mass of solute divided by the volume of the solution, multiplied by 100 %.

i.e  Percent (m/v) =( mass of solute(in gm) / volume of solution(ml) ) × 100 %

Step 2: Extract the data from the question.

Given, Percent (m/V)=26%, Solute=NaCl=14.4 g

Step 3:

Calculation of volume of solution.

By using above formula

Percent (m/v) = (mass of solute/volume of solution) × 100 %

Substituting the values in above equation

26%=(14.4g / volume of solution) ×100%

or, volume of solution = (14.4g / 26%) ×100% = 55.4 ml

hence, the volume of solution = 55.4 ml